OpenStudy (anonymous):

Solve the equation for: two less than four times a number is equal to two more than three times the same number. (2 - 4n =2 +3n)

OpenStudy (anonymous):

help :(

OpenStudy (anonymous):

2-4n=2+3n add 4n both the sides 2-4n+4n=2+3n+4n 2=2+7n 7n=0 n=0

OpenStudy (anonymous):

it can't be zero

OpenStudy (anonymous):

this is wat u've written in ur equation

OpenStudy (anonymous):

can i show you the real question?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

zero is not one of the options

OpenStudy (ghazi):

4n-2=3n+2 this is the question you can subtract 3n on both the sides 4n-3n-2=3n+2-3n n-2=2 add 2 on both the sides n-2+2=2+2 n=4

OpenStudy (anonymous):

yep thats right u wrote rong eqation

OpenStudy (anonymous):

how is it wrong

OpenStudy (ghazi):

because 2 is less than 4 times number not number is less than 2

OpenStudy (anonymous):

okay ghazi

OpenStudy (anonymous):

ur question says 4n-2

OpenStudy (anonymous):

oh my bad, i didn't write that correctly :D

OpenStudy (anonymous):

sorry :D

OpenStudy (anonymous):

yep dats fine (;

OpenStudy (anonymous):

soooo....

OpenStudy (anonymous):

lol watchu wanna do (;

OpenStudy (anonymous):

solve equations

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

i didnt do ny hlp lol

OpenStudy (anonymous):

yes you did

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

@ghazi solved ur problem lol

OpenStudy (anonymous):

okay